# Da Yu: Towards USV-Based Image Captioning for Waterway Surveillance and Scene Understanding

### v1

@@ Abstract
Automated waterway environment perception is crucial for enabling unmanned surface vessels (USVs) to understand their surroundings and make informed decisions. Most existing waterway perception models primarily focus on instance-level object perception paradigms (e.g., detection, segmentation). However, due to the complexity of waterway environments, current perception datasets and models fail to achieve global semantic understanding of waterways, limiting large-scale monitoring and structured log generation. With the advancement of vision-language models (VLMs), we leverage image captioning to introduce WaterCaption, the first captioning dataset specifically designed for waterway environments. WaterCaption focuses on fine-grained, multi-region long-text descriptions, providing a new research direction for visual geo-understanding and spatial scene cognition. Exactly, it includes 20.2k image-text pair data with 1.8 million vocabulary size. Additionally, we propose Da Yu, an edge-deployable multi-modal large language model for USVs, where we propose a novel vision-to-language projector called Nano Transformer Adaptor (NTA). NTA effectively balances computational efficiency with the capacity for both global and fine-grained local modeling of visual features, thereby significantly enhancing the model’s ability to generate long-form textual outputs. Da Yu achieves an optimal balance between performance and efficiency, surpassing state-of-the-art models on WaterCaption and several other captioning benchmarks.
@@ I Introduction
Correspondingly, considering the existence of communication blind spots in waterway environments, traditional cloud-based inference, with an average latency of 300–500 milliseconds over 4G or 5G networks, fails to meet the 100-millisecond response time required by the Inland Collision Avoidance Regulations [19]. To address this limitation and enable real-time generation of waterway monitoring logs, we propose Da Yu [^1] footnotetext: 1Da Yu, is a legendary figure in ancient Chinese history renowned for his extraordinary efforts in controlling floods and protecting waterways. , an edge-deployable MLLM for maritime monitoring. Da Yu fundamentally redefines onboard intelligence by synergistically integrating our novel Nano Transformer Adaptor (NTA), which eliminates the traditional trade-off between computational efficiency and environmental comprehension. Through its dual innovation of Multi-Head Pooling Attention and Dilated Grouped Convolution, the NTA enables linear-complexity processing of high-resolution visual streams while preserving pixel-level fidelity for critical maritime features (e.g., floating debris, vessel identification, and shoreline anomalies). Specifically, the NTA preserves the capability for global context modeling while reducing the parameter count by approximately 24% compared to the convolution-based LDPv2 module [20]. This breakthrough allows Da Yu to perform autonomous, cloud-free situational awareness of dynamic waterways, analysing complex waterway scenarios in real-time on embedded Jetson Orin platforms.
@@ II-A Image Captioning based on Vision-Language Models
Recent advancements in VLMs have significantly improved the performance of image captioning tasks. Models such as CLIP [21], ALBEF [22], BLIP [23], and BLIP-2 [24] have demonstrated remarkable capabilities in aligning visual and textual modalities. CLIP leverages contrastive learning to align image and text embeddings, enabling zero-shot and few-shot learning capabilities across various vision-language tasks. ALBEF introduces an ”align before fuse” mechanism to address challenges in image-text alignment and interaction within the multi-modal domain, achieving notable performance improvements in tasks such as image-text retrieval and cross-modal generation. BLIP and BLIP-2 further enhance this approach by introducing a bootstrapping method to generate synthetic captions and filter noisy ones, achieving state-of-the-art results in tasks like image captioning and visual question answering. The integration of large language models (LLMs) with VLMs has opened new avenues for generating richer and more context-aware captions. For instance, models like LLaVA [25] and Qwen-VL [26] combine the strengths of CLIP and LLMs to handle complex visual-linguistic tasks. However, generating long-text captions for multiple regions within an image remains a significant challenge. Traditional VLMs often struggle with computational efficiency and scalability when handling long-text generation, as they require processing extensive multi-modal interactions. Moreover, these models typically have limited reasoning capabilities compared to multi-modal large language models (MLLMs), making it difficult to maintain coherence and semantic consistency over extended sequences. Another limitation of traditional VLMs is their reliance on object detection models, which primarily focus on foreground objects and often fail to provide precise descriptions of background elements. As noted in recent studies, this limitation arises because detection-based captioning models are trained on datasets with sparse annotations for background regions, leading to incomplete scene descriptions.
@@ II-B Efficient Multi-Modal Large Language Model
However, these models, being pre-trained on general-purpose datasets, still exhibit limited understanding of maritime scenarios. To address this, and to enhance the perceptual capabilities of MLLMs in the maritime domain while fostering an intelligent perception ecosystem for waterborne transportation, we propose Da Yu, a lightweight multimodal large language model designed for maritime perception and optimized for edge inference on USV platforms.
@@ II-C Waterway Perception based on Deep Learning
As a critical component of intelligent urban transportation systems, waterway perception is pivotal to maritime navigation and aquatic environmental conservation. Recent advances in deep learning have led to the widespread adoption of data-driven perception models for waterway monitoring. Prior works primarily focus on specialized tasks: object detection for maritime environments [31, 32, 33], free-space and shoreline segmentation [34, 35, 36], and end-to-end panoptic perception frameworks unifying detection and segmentation [5, 6, 2, 37]. These approaches predominantly rely on multi-sensor fusion for autonomous perception. While recent studies [7, 38] have introduced visual grounding models for intent-aware maritime object localization, the critical task of automated waterway surveillance and log generation, which is essential for navigational safety, remains underexplored due to the lack of dedicated datasets. To bridge this gap, we propose the first multi-regional Waterway Captioning Dataset (WaterCaption), capturing rich semantic descriptions of aquatic environments from USV-deployed RGB cameras in various waterway types. Building upon this, we introduce a novel MLLM called Da Yu specifically optimized for waterway captioning, demonstrating state-of-the-art capabilities in long-form text generation and complex scene understanding under challenging conditions.
@@ III-A Dataset Construction
2) Consider Key Points: In constructing waterway-oriented captions, we systematically consider seven key semantic regions (or focal elements) that comprehensively reflect the functional (navigability), safety-related (risk factors), and ecological (environmental health) attributes of the scene. These seven elements include: (1) waterway objects, which directly impact navigational safety and route planning; (2) waterway buildings, which reveal human interventions and functional demands on the waterway infrastructure; (3) weather conditions, which critically affect image visibility and perception; (4) wave amplitude, which influences the stability of vessel navigation; (5) shore condition, indicating ecological well-being, environmental pollution levels, and the intensity of anthropogenic activities; (6) waterway cleanliness, directly associated with environmental protection and maintenance; and (7) the position of the ego-USV, which ensures the USV remains in a relatively safe spatial relationship with navigation targets and shoreline boundaries. However, not all scenes contain all of the aforementioned seven key elements. In certain waterway scenarios, such as on open lakes or at sea where the shoreline is not visible, the shore condition element is not required to be included in the caption.
@@ IV-B Visual Encoder
We adopt MobileCLIP [41] as the visual encoder, a lightweight and efficient model distilled from the image captioning framework CoCa [42] and OpenAI-CLIP [21]. MobileCLIP inherits CLIP’s capability to capture global, low-frequency, coarse-grained visual label semantics, while simultaneously leveraging the fine-grained, low-frequency visual details learned from CoCa. Compared to SigLIP [43] and the OpenAI-CLIP model, MobileCLIP not only has a smaller parameter footprint but also demonstrates superior image representation performance. Certainly! Here is a professionally written academic-style English translation of your paragraph:
@@ V-B Quantitative Results
Table IX: Comparison of performances with different finetuning methods for Da Yu-L.
@@ V-C Ablation Studies
Figure 12: Comparison between different models on waterway image captioning.

### v2

@@ Abstract
Automated waterway environment perception is crucial for enabling unmanned surface vessels (USVs) to understand their surroundings and make informed decisions. Most existing waterway perception models primarily focus on instance-level object perception paradigms (e.g., detection, segmentation). However, due to the complexity of waterway environments, current perception datasets and models fail to achieve global semantic understanding of waterways, limiting large-scale monitoring and structured log generation. With the advancement of vision-language models (VLMs), we leverage image captioning to introduce WaterCaption, the first captioning dataset specifically designed for waterway environments. WaterCaption focuses on fine-grained, multi-region long-text descriptions, providing a new research direction for visual geo-understanding and spatial scene cognition. Exactly, it includes 20.2k image-text pair data with 1.8 million vocabulary size. Additionally, we propose Da Yu, an edge-deployable multi-modal large language model for USVs, where we propose a novel vision-to-language projector called Nano Transformer Adaptor (NTA). NTA effectively balances computational efficiency with the capacity for both global and fine-grained local modeling of visual features, thereby significantly enhancing the model’s ability to generate long-form textual outputs. Da Yu achieves an optimal balance between performance and efficiency, surpassing state-of-the-art models on WaterCaption and several other captioning benchmarks. The project is available at https://github.com/GuanRunwei/WaterCaption.
@@ I Introduction
Correspondingly, considering the existence of communication blind spots in waterway environments, traditional cloud-based inference, with an average latency of 20 seconds over 4G or 5G networks, fails to meet the 8-second response time required by the Inland Collision Avoidance Regulations [19] for the normal USVs. To address this limitation and enable real-time generation of waterway monitoring logs, we propose Da Yu [^1] footnotetext: 1Da Yu, is a legendary figure in ancient Chinese history renowned for his extraordinary efforts in controlling floods and protecting waterways. , an edge-deployable multi-modal large language model (MLLM) for maritime monitoring. Da Yu fundamentally redefines onboard intelligence by synergistically integrating our novel Nano Transformer Adaptor (NTA), which eliminates the traditional trade-off between computational efficiency and environmental comprehension. Through its dual innovation of Multi-Head Pooling Attention and Dilated Grouped Convolution, the NTA enables linear-complexity processing of high-resolution visual streams while preserving pixel-level fidelity for critical maritime features (e.g., floating debris, vessel identification, and shoreline anomalies). Specifically, the NTA preserves the capability for global context modeling while reducing the parameter count by approximately 24% compared to the convolution-based LDPv2 module [20]. This breakthrough allows Da Yu to perform autonomous, cloud-free situational awareness of dynamic waterways, analysing complex waterway scenarios in real-time on embedded Jetson Orin platforms.
@@ II-A Image Captioning based on Vision-Language Models
Recent advancements in VLMs have significantly improved the performance of image captioning tasks. Models such as CLIP [21], ALBEF [22], BLIP [23], and BLIP-2 [24] have demonstrated remarkable capabilities in aligning visual and textual modalities. CLIP leverages contrastive learning to align image and text embeddings, enabling zero-shot and few-shot learning capabilities across various vision-language tasks. ALBEF introduces an ”align before fuse” mechanism to address challenges in image-text alignment and interaction within the multi-modal domain, achieving notable performance improvements in tasks such as image-text retrieval and cross-modal generation. BLIP and BLIP-2 further enhance this approach by introducing a bootstrapping method to generate synthetic captions and filter noisy ones, achieving state-of-the-art results in tasks like image captioning and visual question answering. The integration of large language models (LLMs) with VLMs has opened new avenues for generating richer and more context-aware captions. For instance, models like LLaVA [25] and Qwen-VL [26] combine the strengths of CLIP and LLMs to handle complex visual-linguistic tasks. However, generating long-text captions for multiple regions within an image remains a significant challenge. Traditional VLMs often struggle with computational efficiency and scalability when handling long-text generation, as they require processing extensive multi-modal interactions. Moreover, these models typically have limited reasoning capabilities compared to multi-modal large language models (MLLMs), making it difficult to maintain coherence and semantic consistency over extended sequences. Another limitation of traditional VLMs is their reliance on object detection models, which primarily focus on foreground objects and often fail to provide precise descriptions of background elements [27, 28, 29]. As noted in recent studies, this limitation arises because detection-based captioning models are trained on datasets with sparse annotations for background regions [30, 31], leading to incomplete scene descriptions [32, 33].
@@ II-B Efficient Multi-Modal Large Language Model
However, these models, being pre-trained on general-purpose datasets, still exhibit limited understanding of maritime scenarios. To address this, and to enhance the perceptual capabilities of MLLMs in the maritime domain while fostering an intelligent perception ecosystem for waterborne transportation, we propose Da Yu, a lightweight MLLM designed for maritime perception and optimized for edge inference on USV platforms.
@@ II-C Waterway Perception based on Deep Learning
As a critical component of intelligent urban transportation systems, waterway perception is pivotal to maritime navigation and aquatic environmental conservation. Recent advances in deep learning have led to the widespread adoption of data-driven perception models for waterway monitoring. Prior works primarily focus on specialized tasks: object detection for maritime environments [5], free-space and shoreline segmentation [38], and end-to-end panoptic perception frameworks unifying detection and segmentation [2]. These approaches predominantly rely on multi-sensor fusion for autonomous perception. While recent studies [7, 39] have introduced visual grounding models for intent-aware maritime object localization, the critical task of automated waterway surveillance and log generation, which is essential for navigational safety, remains underexplored due to the lack of dedicated datasets. To bridge this gap, we propose the first multi-regional Waterway Captioning Dataset (WaterCaption), capturing rich semantic descriptions of aquatic environments from USV-deployed RGB cameras in various waterway types. Building upon this, we introduce a novel MLLM called Da Yu specifically optimized for waterway captioning, demonstrating state-of-the-art capabilities in long-form text generation and complex scene understanding under challenging conditions.
@@ III-A Dataset Construction
2) Consider Key Points: In constructing waterway-oriented captions, we systematically consider seven key semantic regions (or focal elements) that comprehensively reflect the functional (navigability), safety-related (risk factors), and ecological (environmental health) attributes of the scene. These seven elements include: (1) waterway objects, which directly impact navigational safety and route planning in the channel; (2) waterway buildings, which reveal human interventions and functional demands on the waterway infrastructure; (3) weather conditions, which critically affect image visibility and perception; (4) wave amplitude, which influences the stability of vessel navigation; (5) shore condition, indicating ecological well-being, environmental pollution levels, and the intensity of anthropogenic activities; (6) waterway cleanliness, directly associated with environmental protection and maintenance; and (7) the position of the ego-USV, which ensures the USV remains in a relatively safe spatial relationship with navigation targets and shoreline boundaries. However, not all scenes contain all of the aforementioned seven key elements. In certain waterway scenarios, such as on open lakes or at sea where the shoreline is not visible, the shore condition element is not required to be included in the caption.
@@ IV-B Visual Encoder
We adopt MobileCLIP [42] as the visual encoder, a lightweight and efficient model distilled from the image captioning framework CoCa [43] and OpenAI-CLIP [21]. MobileCLIP inherits CLIP’s capability to capture global, low-frequency, coarse-grained visual label semantics, while simultaneously leveraging the fine-grained, low-frequency visual details learned from CoCa. Compared to SigLIP [44] and the OpenAI-CLIP model, MobileCLIP not only has a smaller parameter footprint but also demonstrates superior image representation performance.
@@ V-B Quantitative Results
Table IX: Comparison of different finetuning methods for Da Yu-L.
@@ V-C Ablation Studies
Figure 12: Comparison between different models on waterway image captioning for different key elements.